You buy a house of $450,000 today. You put a down payment of 20% and borrow a fixed-rate mortgage of $360,000 with interest rate of 4% and 15 years. After 3 years, your house is appreciated to the value of $550,000 and market interest rate goes up to 6.5%. How much money will you make in book after 3 years?  
Answer #1

Loan outstanding after 3 years=450000*(1-20%)*(1+4%/12)^(12*3)-(450000*(1-20%)*(4%/12)/(1-1/(1+4%/12)^(12*15)))/(4%/12)*((1+4%/12)^(12*3)-1)=$304,145.09

Equity after 3 years=550000-304145.09=$245,854.91

Equity iniitally=450000*20%=$90,000.00

Money made in book=245854.91-90000.00=$155,854.91